THE MAGELLANIC CLOUDS NEWSLETTER

An electronic publication dedicated to the Magellanic Clouds, and astrophysical phenomena therein

No. 96 — 1 December 2008

http://www.astro.keele.ac.uk/MCnews

Editor: Jacco van Loon

Editorial

Dear Colleagues,

It is my pleasure to present you the 96^{th} issue of the Magellanic Clouds Newsletter.

It is full of interesting new results, for instance on interpreting the latest determinations of the motions of the Clouds, on the transition from an atomic to a molecular interstellar medium, the discovery of multiple stellar populations in many populous star clusters in the Magellanic Clouds, and results from Spitzer and AKARI on star formation, dust in the SMC tail, variable stars including the mass loss from Cepheids, and on supernova remnants. Oh yes, and three more articles on SN(R) 1987A alone!

If you are looking for a postdoctoral position, there is one open at the Observatory of the Côte d’Azur, in Southern France. A nice place, and an exciting project to work on.

The next issue will be distributed on the 1^{st} of February 2009; the deadline for contributions is the 31^{st} of January. With best wishes for the festive season and a Happy New Year,

Editorially Yours, Jacco van Loon

Refereed Journal Papers

Dynamical Masses for the Large Magellanic Cloud Massive Binary System [L72] LH 54-425

S.J. Williams^{1}, D.R. Gies^{1}, T.J. Henry^{1}, J.A. Orosz^{2}, M.V. McSwain^{3}, T.C. Hillwig^{4}, L.R. Penny^{5}, G. Sonneborn^{6}, R. Iping^{7,8}, K.A. van der Hucht^{9,10} and L. Kaper^{10}

1Center for High Angular Resolution Astronomy and Department of Physics and Astronomy, Georgia State University, P.O. Box 4106, Atlanta, GA 30302-4106, USA 2Department of Astronomy, San Diego State University, 5500 Campanile Drive, San Diego, CA 92182-1221, USA 3Department of Physics, Lehigh University, 16 Memorial Drive East, Bethlehem, PA 18015, USA 4Department of Physics and Astronomy, Valparaiso University, Valparaiso, IN 46383, USA 5Department of Physics and Astronomy, College of Charleston, 101 Science Center, 58 Coming Street, Charleston, SC 29424, USA 6NASA Goddard Space Flight Center, Code 681, Greenbelt, MD 20771, USA 7Department of Physics, Catholic University of America, Washington, DC, USA 8NASA Goddard Space Flight Center, Code 665, Greenbelt, MD 20771, USA 9SRON Netherlands Institute for Space Research, Sorbonnelaan 2, 3584CA Utrecht, The Netherlands 10Astronomical Institute Anton Pannekoek, University of Amsterdam, Kruislaan 403, 1098 SJ Amsterdam, The Netherlands

We present results from an optical spectroscopic investigation of the massive binary system [L72] LH 54-425 in the LH 54 OB association in the Large Magellanic Cloud. We revise the ephemeris of [L72] LH 54-425 and find an orbital period of 2.24741 ± 0.00004 days. We find spectral types of O3 V for the primary and O5 V for the secondary. We made a combined solution of the radial velocities and previously published V -band photometry to determine the inclination for two system configurations, i = 52_{−3}^{+2} degrees for the configuration of the secondary star being more tidally distorted and i = 55^{∘} ± 1^{∘} for the primary as the more tidally distorted star. We argue that the latter case is more probable, and this solution yields masses and radii of M_{1} = 47 ± 2 M_{⊙} and R_{1} = 11.4 ± 0.1 R_{⊙} for the primary, and M_{2} = 28 ± 1 M_{⊙} and R_{2} = 8.1 ± 0.1 R_{⊙} for the secondary. Our analysis places LH 54-425 amongst the most massive stars known. Based on the position of the two stars plotted on a theoretical HR diagram, we find the age of the system to be ∼1.5 Myr.

Published in Astrophysical Journal, 682, 492 (2008)

Improved analysis of SN 1987A antineutrino events

Giulia Pagliaroli^{1,2}, Francesco Vissani^{1}, Maria Laura Costantini^{1,2} and Aldo Ianni^{1}

1INFN, Laboratori Nazionali del Gran Sasso, Italy 2University of L’Aquila, Italy

We propose a new parameterization of the antineutrino flux from core collapse supernovae, that allows an interpretation of its astrophysical parameters within the Bethe and Wilson scenario for the explosion, and that leads to a reasonable (smooth) behavior of the average energy and of the luminosity curve. We apply it to analyze the events observed by Kamiokande-II, IMB and Baksan detectors in correlation with SN 1987A. For the first time, we consider in the same analysis all data characteristics: times, energies and angles of the observed events. We account for the presence of background and evaluate the impact of neutrino oscillations. The hypothesis that the initial luminous phase of emission (accretion) is absent can be rejected at the 2% significance level. Without the need to impose external priors in the likelihood analysis, the best-fit values of the astrophysical parameters are found to be in remarkable agreement with the expectations of the standard core-collapse scenario; in particular, the electron antineutrino-sphere radius is 16 km, the duration of the accretion phase is found to be 0.55 s, and the initial accreting mass is 0.22 M_{⊙}. Similarly the total energy emitted in neutrinos is 2.2 × 10^{53} erg, again close to the expectations. The errors on the parameters are evaluated and found to be relatively large, consistently with the limited number of detected events; the two dimensional confidence regions, that demonstrate the main correlations between the parameters, are also given.

Submitted to Physical Review D Available from arXiv:0810.0466

Spatial motion of the Magellanic Clouds. Tidal models ruled out?

Adam Růžička^{1,2}, Christian Theis^{2} and Jan Palouš^{1}

1Astronomical Institute, Academy of Sciences of the Czech Republic, v.v.i., Boční II 1401, 141 31, Prague, Czech Republic 2Institut für Astronomie der Universität Wien, Türkenschanzstraße 17, A-1180 Wien, Austria

Recently, Kallivayalil et al. derived new values of the proper motion for the Large and Small Magellanic Clouds (LMC and SMC, respectively). The spatial velocities of both Clouds are unexpectedly higher than their previous values resulting from agreement between the available theoretical models of the Magellanic System and the observations of neutral hydrogen (H i) associated with the LMC and the SMC. Such proper motion estimates are likely to be at odds with the scenarios for creation of the large-scale structures in the Magellanic System suggested so far. We investigated this hypothesis for the pure tidal models, as they were the first ones devised to explain the evolution of the Magellanic System, and the tidal stripping is intrinsically involved in every model assuming the gravitational interaction. The parameter space for the Milky Way (MW)–LMC–SMC interaction was analyzed by a robust search algorithm (genetic algorithm) combined with a fast restricted N-body model of the interaction. Our method extended the known variety of evolutionary scenarios satisfying the observed kinematics and morphology of the Magellanic large-scale structures. Nevertheless, assuming the tidal interaction, no satisfactory reproduction of the H i data available for the Magellanic Clouds was achieved with the new proper motions. We conclude that for the proper motion data by Kallivayalil et al., within their 1-σ errors, the dynamical evolution of the Magellanic System with the currently accepted total mass of the MW cannot be explained in the framework of pure tidal models. The optimal value for the western component of the LMC proper motion was found to be μ_{lmc}^{W}> ∼ − 1.3 mas yr^{−1} in case of tidal models. It corresponds to the reduction of the Kallivayalil et al. value for μ_{lmc}^{W} by ≈ 40% in its magnitude.

Accepted for publication in ApJ Available from arXiv:0810.0968

Multiple stellar populations in Magellanic Clouds clusters. I. An ordinary feature for intermediate age globulars in the LMC?

A.P. Milone^{1}, L.R. Bedin^{2}, G. Piotto^{1} and J. Anderson^{2}

1Dipartimento di Astronomia, Università di Padova, Vicolo dell’Osservatorio 3, Padova, I-35122, Italy 2Space Telescope Science Institute, 3700 San Martin Drive, Baltimore, MD 21218, USA

The discovery of multiple main sequences (MS) in the massive clusters NGC 2808 and ω Centauri, and multiple subgiant branches in NGC 1851 and NGC 6388 has challenged the long-held paradigm that globular clusters consist of simple stellar populations. This evolving picture has been further complicated by recent photometric studies of the Large Magellanic Cloud (LMC) intermediate-age clusters, where the main sequence turn-off (MSTO) was found to be bimodal (NGC 1806 and NGC 1846) or broadened (NGC 1783 and NGC 2173). We have undertaken a study of archival HST images of Large and Small Magellanic Cloud clusters with the aim of measuring the frequency of clusters with evidence of multiple or prolonged star formation events and determining their main properties. We found useful images for 53 clusters that cover a wide range of ages. In this paper, we analyse the Color-Magnitude Diagrams (CMD) of sixteen intermediate-age (∼1–3 Gyr) LMC clusters. The data were reduced by using the method developed by Anderson et al. (2008) and the photometry has been corrected for differential reddening (where required). We find that eleven clusters show an anomalous spread (or split) in color and magnitude around the MSTO, even though the other main features of the CMD (MS, red giant branch, asymptotic giant branch) are narrow and the horizontal branch (HB) red-clump well defined. By using the CMD of the stars in regions that surround the cluster, we demonstrate that the observed feature is unequivocally associated to the clusters. We use artificial-star tests to demonstrate that the spread (or split) is not an artifact due to photometric errors or binaries. We confirm that two clusters (NGC 1806 and NGC 1846) clearly exhibit two distinct MSTOs and observe, for the first time, a double MSTO in NGC 1751. In these three clusters the population corresponding to the brighter MSTO includes more than two-thirds of cluster stellar population. We confirm the presence of multiple stellar populations in NGC 1783. Our photometry strongly suggests that the MSTO of this cluster is formed by two distinct branches. In seven clusters (ESO 057-SC 075, HODGE 7, NGC 1852, NGC 1917, NGC 1987, NGC 2108, and NGC 2154) we observed an intrinsic broadening of the MSTO that may suggest that these clusters have experienced a prolonged period of star formation that span a period between 150 and 250 Myr. The CMDs of IC 2146, NGC 1644, NGC 1652, NGC 1795 and NGC 1978 show no evidence of spread or bimodality within our photometric precision. In summary 70±25% of our sample are not consistent with the simple, single stellar population hypothesis.

Accepted for publication in A&A Available from arXiv:0810.2558 and from ftp://ftp.stsci.edu/outside-access/out.going/bedin/LMC/

Radial velocities, dynamics of stars and nebulosities with GAIA and VLT-GIRAFFE

C. Martayan^{1,2}, Y. Frémat^{1}, R. Blomme^{1}, A. Jonckheere^{1}, C. Delle-Luche^{2}, P. Sartoretti^{2}, D. Katz^{2}, Y. Viala^{2}, M. Floquet^{2}, A.-M. Hubert^{2} and C. Neiner^{2}

1Royal Observatory of Belgium, 3 Avenue circulaire, 1180 Brussels, Belgium 2GEPI, Observatoire de Paris, CNRS, Université Paris Diderot, 5 place Jules Janssen, 92195 Meudon Cedex, France

This document is divided in two parts. The first part deals with the radial velocities (RV) distributions for B-type stars and nebulosities observed with the VLT-GIRAFFE in the Large and Small Magellanic Clouds towards the open clusters NGC 2004 and NGC 330. Thanks to the resolution of GIRAFFE spectra, we found that the RV distribution for the nebulosities in the LMC is bi-modal. This bi-modality can be interpreted, in term of dynamics, by the expansion of the LMC 4 superbubble. The second part deals with the GAIA space mission and the determination of the radial velocities by using Radial Velocity Spectrometer (RVS) spectra. The methods to determine the radial velocities are presented as well as preliminary results on simulated RVS spectra.

Accepted for publication in AN Available from arXiv:0809.3241

The Physical Properties and Effective Temperature Scale of O-type

Stars as a Function of Metallicity. III. More Results from the Magellanic Clouds

Philip Massey^{1}, Amanda M. Zangari^{1,2}, Nidia I. Morrell^{3}, Joachim Puls^{4}, Kathleen DeGioia-Eastwood^{5}, Fabio Bresolin^{6} and Rolf-Peter Kudritzki^{6}

1Lowell Observatory, USA 2MIT, USA 3Las Campanas Observatory, Chile 4Universitäts-Sternwarte München, Germany 5Northern Arizona University, USA 6Institute for Astronomy, University of Hawai’i, USA

In order to better determine the physical properties of hot, massive stars as a function of metallicity, we obtained very high SNR optical spectra of 26 O and early B stars in the Magellanic Clouds. These allow accurate modeling even in cases where the He i λ4471 line has an equivalent width of only a few tens of mÅ. The spectra were modeled with FASTWIND, with good fits obtained for 18 stars; the remainder show signatures of being binaries. We include stars in common to recent studies to investigate possible systematic differences. The “automatic” FASTWIND modeling method of Mokiem and collaborators produced temperatures 1100 K hotter on the average, presumably due to the different emphasis given to various temperature-sensitive lines. More significant, however, is that the automatic method always produced some “best” answer, even for stars we identify as composite (binaries). The temperatures found by the TLUSTY/CMFGEN modeling of Bouret, Heap, and collaborators yielded temperatures 1000 K cooler than ours, on average. Significant outliers were due either to real differences in the data (some of the Bouret/Heap data were contaminated by moonlight continua) or the fact we could detect the He i line needed to better constrain the temperature. Our new data agrees well with the effective temperature scale we presented previously. We confirm that the “Of” emission lines do not track luminosity classes in the exact same manner as in Milky Way stars. We revisit the the issue of the “mass discrepancy”, finding that some of the stars in our sample do have spectroscopic masses that are significantly smaller than those derived from stellar evolutionary models. We do not find that the size of the mass discrepancy is simply related to either effective temperature or surface gravity.

Accepted for publication in ApJ Available from arXiv:0810.3962 and from http://www.lowell.edu/users/massey/o2uv.pdf.gz

Testing Mass Loss in Large Magellanic Cloud Cepheids using Infrared and Optical Observations

Hilding R. Neilson^{1}, Chow-Choong Ngeow^{2}, Shashi M. Kanbur^{3} and John B. Lester^{4}

1University of Toronto, Canada 2UIUC, USA 3SUNY Oswego, USA 4University of Toronto Mississauga, Canada

It has been claimed that Period-Luminosity relations derived from infrared observations of Large Magellanic Cloud (LMC) Cepheids are less dependent on the metallicity of the Cepheids. In this work, infrared observations of LMC Cepheids from the SAGE survey are combined with OGLE II optical observation to model and predict mass-loss rates. The mass-loss rates are fit to the data and are predicted to range from about 10^{−12} to 10^{−7} M_{⊙} yr^{−1}; however, the rates depend on the assumed value of the dust-to-gas ratio. By comparing the relations derived from observations to the relations derived from predicted infrared stellar luminosities from the mass-loss model, it is shown mass loss affects the structure and scatter of the infrared Period–Luminosity relation. Mass loss produces shallower slopes of the infrared relations and a lower zero point. There is also evidence for non-linearity in the predicted Period–Luminosity relations, and it is argued that mass loss produces larger infrared excess at lower periods, which affects the slope and zero point, making the PL relations more linear in the wavelength range of 3.6 to 5.8 μm. Because the dust-to-gas ratio is metallicity dependent and mass loss may have a metallicity dependence, infrared Period–Luminosity relations have additional uncertainty due to metallicity.

Accepted for publication in ApJ Available from arXiv:0810.3001

The Spatial Evolution of Stellar Structures in the LMC

Nate Bastian^{1,2}, Mark Gieles^{3}, Barbara Ercolano^{1,4} and Robert Gutermuth^{4}

1IoA - Cambridge, UK 2UCL, UK 3ESO - Santiago, Chile 4CfA, USA

We present an analysis of the spatial distribution of various stellar populations within the Large Magellanic Cloud.

We combine mid-infrared selected young stellar objects, optically selected samples with mean ages between ∼ 9 and ∼ 1000, and existing stellar cluster catalogues to investigate how stellar structures form and evolve within the LMC. For the analysis we use Fractured Minimum Spanning Trees, the statistical Q parameter, and the two-point correlation function. Restricting our analysis to young massive (OB) stars we confirm our results obtained for M 33, namely that the luminosity function of the groups is well described by a power-law with index −2, and that there is no characteristic length-scale of star-forming regions. We find that stars in the LMC are born with a large amount of substructure, consistent with a 2D fractal distribution with dimension ∼ 1.8 and evolve towards a uniform distribution on a timescale of ∼ 175 Myr. This is comparable to the crossing time of the galaxy and we suggest that stellar structure, regardless of spatial scale, will be eliminated in a crossing time. This may explain the smooth distribution of stars in massive/dense young clusters in the Galaxy, while other, less massive, clusters still display large amounts of structure at similar ages. By comparing the stellar and star cluster distributions and evolving timescales, we show that infant mortality of clusters (or ’popping clusters’) have a negligible influence on galactic structure. Finally, we quantify the influence of the elongation, differential extinction, and contamination of a population on the measured Q value.

Accepted for publication in MNRAS Available from arXiv:0810.3190

High-Resolution X-ray Spectroscopy of SNR 1987A: Chandra LETG and HETG Observations in 2007

S.A. Zhekov^{1}, R. McCray^{1}, D. Dewey^{2}, C.R. Canizares^{2}, K.J. Borkowski^{3}, D.N. Burrows^{4} and S. Park^{4}

1JILA, University of Colorado, Boulder, USA 2MIT, Kavli Institute, Cambridge, USA 3NCSU, Raleigh, USA 4Pennsylvania State University, State College, USA

We present an extended analysis of the deep Chandra LETG and HETG observations of the supernova remnant 1987A (SNR 1987A) carried out in 2007. The global fits to the grating spectra show that the temperature of the X-ray emitting plasma in the slower shocks in this system has remained stable for the last three years, while that in the faster shocks has decreased. This temperature evolution is confirmed by the first light curves of strong X-ray emission lines and their ratios. On the other hand, bulk gas velocities inferred from the X-ray line profiles are too low to account for the post-shock plasma temperatures inferred from spectral fits. This suggests that the X-ray emission comes from gas that has been shocked twice, first by the blast wave and again by shocks reflected from the inner ring of SNR 1987A. A new model that takes these considerations into account gives support to this physical picture.

Accepted for publication in Astrophysical Journal Available from arXiv:8010.5313

Time evolution of the line emission from the inner circumstellar ring of SN 1987A and its hot spots

Per Gröningsson^{1}, Claes Fransson^{1}, Bruno Leibundgut^{2}, Peter Lundqvist^{1}, Peter Challis^{3}, Roger A. Chevalier^{4} and Jason Spyromilio^{2}

1Stockholm Observatory, Stockholm University, AlbaNova University Center, SE-106 91 Stockholm, Sweden 2European Southern Observatory, Karl-Schwarzschild-Straße 2, D-85748 Garching, Germany 3Harvard-Smithsonian Center for Astrophysics, 60 Garden Street, MS-19, Cambridge, MA 02138, U.S.A. 4Department of Astronomy, University of Virginia, P.O. Box 400325, Charlottesville, VA 22904, U.S.A.

We present seven epochs between October 1999 and November 2007 of high resolution VLT/UVES echelle spectra of the ejecta-ring collision of SN 1987A. The fluxes of most of the narrow lines from the unshocked gas decreased by a factor of 2–3 during this period, consistent with the decay from the initial ionization by the shock break-out. However, [O iii] in particular shows an increase up to day ∼ 6800. This agrees with radiative shock models where the pre-shocked gas is heated by the soft X-rays from the shock. The evolution of the [O iii] line ratio shows a decreasing temperature of the unshocked ring gas, consistent with a transition from a hot, low density component which was heated by the initial flash ionization to the lower temperature in the pre-ionized gas ahead of the shocks. The line emission from the shocked gas increases rapidly as the shock sweeps up more gas. We find that the neutral and high ionization lines follow the evolution of the Balmer lines roughly, while the intermediate ionization lines evolve less rapidly. Up to day ∼ 6800, the optical light curves have a similar evolution to that of the soft X-rays. The break between day 6500 and day 7000 for [O iii] and [Ne iii] is likely due to recombination to lower ionization levels. Nevertheless, the evolution of the [Fe xiv] line, as well as the lines from the lowest ionization stages, continue to follow that of the soft X-rays, as expected. There is a clear difference in the line profiles between the low and intermediate ionization lines, and those from the coronal lines at the earlier epochs. This shows that these lines arise from regions with different physical conditions, with at least a fraction of the coronal lines coming from adiabatic shocks. At later epochs the line widths of the low ionization lines, however, increase and approach those of the high ionization lines of [Fe x–xiv]. The Hα line profile can be traced up to ∼ 500 km s^{−1} at the latest epoch. This is consistent with the cooling time of shocks propagating into a density of (1 − 4) × 10^{4} cm^{−3}. This means that these shocks are among the highest velocity radiative shocks observed.

Accepted for publication in Astronomy & Astrophysics Available from arXiv:0810.2661

The Atomic to Molecular Transition in Galaxies. II: H i and H_{2} Column Densities

Mark R. Krumholz^{1,2}, Christopher F. McKee^{3} and Jason Tumlinson^{4}

1Hubble Fellow, Department of Astrophysical Sciences, Princeton University, Peyton Hall, Princeton, NJ 08544, USA 2Department of Astronomy & Astrophysics, University of California, Santa Cruz, Interdisciplinary Sciences Building, Santa Cruz, CA 95064, USA 3Departments of Physics and Astronomy, University of California, Berkeley, Campbell Hall, Berkeley, CA 94720-7304, USA 4Space Telescope Science Institute, 3700 San Martin Dr., Baltimore, MD 21218, USA

Gas in galactic disks is collected by gravitational instabilities into giant atomic-molecular complexes, but only the inner, molecular parts of these structures are able to collapse to form stars. Determining what controls the ratio of atomic to molecular hydrogen in complexes is therefore a significant problem in star formation and galactic evolution. In this paper we use the model of H_{2} formation, dissociation, and shielding developed in the previous paper in this series to make theoretical predictions for atomic to molecular ratios as a function of galactic properties. We find that the molecular fraction in a galaxy is determined primarily by its column density and secondarily by its metallicity, and is to good approximation independent of the strength of the interstellar radiation field. We show that the column of atomic hydrogen required to shield a molecular region against dissociation is ∼ 10 M_{⊙} pc^{−2} at solar metallicity. We compare our model to data from recent surveys of the Milky Way and of nearby galaxies, including the Magellanic Clouds, and show that the both the primary dependence of molecular fraction on column density and the secondary dependence on metallicity that we predict are in good agreement with observed galaxy properties.

Accepted for publication in ApJ Available from arXiv:0811.0004 and from http://www.ucolick.org/∼krumholz/publications.html

The 6-GHz Multibeam Maser Survey I. Techniques

J.A. Green^{1,2}, J.L. Caswell^{2}, G.A. Fuller^{1}, A. Avison^{1}, S.L. Breen^{2,3}, K. Brooks^{2}, M.G. Burton^{4}, A.

Chrysostomou^{5}, J. Cox^{6}, P.J. Diamond^{1}, S.P. Ellingsen^{3}, M.D. Gray^{1}, M.G. Hoare^{7}, M.R.W. Masheder^{8}, N.M. McClure-Griffiths^{2}, M. Pestalozzi^{5,9}, C. Phillips^{2}, L. Quinn^{1}, M.A. Thompson^{5}, M.A. Voronkov^{2}, A. Walsh^{10}, D.

Ward-Thompson^{6}, D. Wong-McSweeney^{1}, J.A. Yates^{11} and R.J. Cohen^{1}

1Jodrell Bank Centre for Astrophysics, UK 2Australia Telescope National Facility, Australia 3University of Tasmania, Australia 4University of New South Wales, Australia 5University of Hertfordshire, UK 6Cardiff University, UK 7University of Leeds, UK 8Bristol University, UK 9Göteborgs Universitet, Sweden 10James Cook University, Australia 11University College London, UK

A new 7-beam 6–7 GHz receiver has been built to survey the Galaxy and the Magellanic Clouds for newly forming highmass stars that are pinpointed by strong methanol maser emission at 6668 MHz. The receiver was jointly constructed by Jodrell Bank Observatory (JBO) and the Australia Telescope National Facility (ATNF) and allows simultaneous coverage at 6668 and 6035 MHz. It was successfully commissioned at Parkes in January 2006 and is now being used to conduct the Parkes-Jodrell multibeam maser survey of the Milky Way. This will be the first systematic survey of the entire Galactic plane for masers of not only 6668-MHz methanol, but also 6035-MHz excited-state hydroxyl. The survey is two orders of magnitude faster than most previous systematic surveys and has an rms noise level of ∼ 0.17 Jy. This paper describes the observational strategy, techniques and reduction procedures of the Galactic and Magellanic Cloud surveys, together with deeper, pointed, follow-up observations and complementary observations with other instruments. It also includes an estimate of the survey detection efficiency. The 111 days of observations with the Parkes telescope have so far yielded > 800 methanol sources, of which ∼ 350 are new discoveries. The whole project will provide the first comprehensive Galaxy-wide catalogue of 6668-MHz and 6035-MHz masers.

Accepted for publication in MNRAS Available from arXiv:0810.5201

Period–Luminosity Relations Derived from the OGLE-III Fundamental Mode Cepheids

C. Ngeow^{1}, S. Kanbur^{2}, H. Neilson^{3}, A. Nanthakumar^{2} and J. Buonaccorsi^{4}

1University of Illinois (Urbana-Champaign), USA 2SUNY-Oswego, USA 3University of Toronto, Canada 4University of Massachusetts, USA

In this Paper, we have derived Cepheid period-luminosity (P–L) relations for the Large Magellanic Cloud (LMC) fundamental mode Cepheids, based on the data released from OGLE-III. We have applied an extinction map to correct for the extinction of these Cepheids. In addition to the V IW band P–L relations, we also include JHK and four Spitzer IRAC band P–L relations, derived by matching the OGLE-III Cepheids to the 2MASS and SAGE datasets, respectively. We also test the non-linearity of the Cepheid P–L relations based on extinction-corrected data. Our results (again) show that the LMC P–L relations are non-linear in V IJH bands and linear in KW and the four IRAC bands, respectively.

Accepted for publication in Astrophysical Journal Available from arXiv:0811.2000

Supernova Remnants in the AKARI IRC Survey of the Large Magellanic Cloud

Ji Yeon Seok ^{1}, Bon-Chul Koo^{1}, Takashi Onaka^{2}, Yoshifusa Ita^{3}, Ho-Gyu Lee^{1}, Jae-Joon Lee^{4}, Dae-Sik Moon^{5}, Itsuki Sakon^{2}, Hidehiro Kaneda^{6}, Hyung Mok Lee^{1}, Myung Gyoon Lee^{1} and Sung Eun Kim^{7}

1Seoul National University, Korea 2University of Tokyo, Japan 3National Astronomical Observatory of Japan, Japan 4Pennsylvania State University, USA 5University of Toronto, Canada 6Institute of Space and Astronautical Science, Japan Aerospace Exploration Agency, Japan 7Sejong University, Korea

We present the near- to mid-infared study of supernova remnants (SNRs) using the AKARI IRC Survey of the Large Magellanic Cloud (LMC). The LMC survey observed about a 10 square degree area of the LMC in five bands centered at 3, 7, 11, 15, and 24 μm using the Infrared Camera (IRC) aboard AKARI. The number of SNRs in the survey area is 21, which is about a half of the known LMC SNRs. We systematically examined the AKARI images and identified eight SNRs with distinguishable infrared emission. All of them were detected at > ∼10 μm and some at 3 and 7 μm, too. We present their AKARI images and fluxes. In the 11/15 μm versus 15/24 μm color–color diagram, the SNRs appear to be aligned along a modified blackbody curve, representing thermal emission from dust at temperatures between 90 and 190 K. There is a good correlation between the 24 μm and X-ray fluxes of the SNRs. It was also found that there is a good correlation between the 24 μm and radio fluxes even if there is no direct physical connection between them. We considered the origin of the detected mid-infrared emission in individual SNRs. We conclude that the mid-infrared emissions in five SNRs that show morphologies similar to the X-rays are dominated by thermal emission from hot dust heated by X-ray emitting plasma. Their 15/24 μm color temperatures are generally higher than the Spitzer 24/70 μm color temperatures, which suggests that a single-temperature dust model cannot describe the full spectral energy distribution (SED) of the SNRs. It also implies that our understanding of the full SED is essential for estimating the dust destruction rate of grains by SNR shocks.

Accepted for publication in PASJ Available from arXiv:0811.1403

Variable Evolved Stars and YSOs Discovered in the Large Magellanic Cloud using the SAGE Survey

Um P. Vijh^{1,2}, M. Meixner^{2}, B. Babler ^{3}, M. Block^{4}, S. Bracker^{3}, C. W. Engelbracht^{4}, B. For^{4}, K. Gordon^{2}, J. Hora^{5}, R. Indebetouw^{6}, C. Leitherer^{2}, M. Meade^{3}, K. Misselt^{4}, M. Sewilo^{2}, S. Srinivasan^{7} and B. Whitney^{8}

1Present Address: Rittter Astrophysical Research Center, University of Toledo, Toledo, OH 43606, USA 2Space Telescope Science Institute, Baltimore, MD 21218, USA 3University of Wisconsin, Madison, WI 53706, USA 4University of Arizona, Tucson, AZ 85719, USA 5CfA/Harvard, Cambridge, MA 02138, USA 6University of Virginia, Charlottesville, VA 22903, USA 7Johns Hopkins University, Baltimore, MD 21218, USA 8Space Science Institute, Madison, WI 53716, USA

We present initial results and source lists of variable sources in the Large Magellanic Cloud (LMC) for which we detect thermal infrared variability from the SAGE (Surveying the Agents of a Galaxy’s Evolution) survey, which had 2 epochs of photometry separated by three months. The SAGE survey mapped a 7^{∘} × 7^{∘} region of the LMC using the IRAC and the MIPS instruments on board Spitzer. Variable sources are identified using a combination of the IRAC 3.6, 4.5, 5.8, 8.0 μm bands and the MIPS 24 μm bands. An error-weighted flux difference between the two epochs is used to assess the variability. Of the ∼ 3 million sources detected at both epochs we find ∼ 2, 000 variable sources for which we provide electronic catalogs. Most of the variable sources can be classified as asymptotic giant branch (AGB) stars. A large fraction (> 66%) of the extreme AGB stars are variable and only smaller fractions of carbon-rich (6.1%) and oxygen-rich (2.0%) stars are detected as variable. We also detect a population of variable young stellar object candidates.

Accepted for publication in The Astronomical Journal Available from arXiv:0811.0408

The Dust-to-Gas Ratio in the Small Magellanic Cloud Tail

Karl Gordon^{1}, C. Bot, E. Muller, K.A. Misselt, A. Bolatto, J.-P. Bernard, W. Reach, C. W. Engelbracht, B. Babler, S. Bracker, M. Block, G.C. Clayton, J. Hora, R. Indebetouw, F.P. Israel, A. Li, S. Madden, M. Meade, M. Meixner,

M. Sewilo, B. Shiao, L.J. Smith, J.Th. van Loon and B.A. Whitney

1Space Telescope Science Institute, USA

The Tail region of the Small Magellanic Cloud (SMC) was imaged using the MIPS instrument on the Spitzer Space Telescope as part of the SAGE-SMC Spitzer Legacy. Diffuse infrared emission from dust was detected in all the MIPS bands. The Tail gas-to-dust ratio was measured to be 1200 ± 350 using the MIPS observations combined with existing IRAS and H i observations. This gas-to-dust ratio is higher than the expected 500–800 from the known Tail metallicity indicating possible destruction of dust grains. Two cluster regions in the Tail were resolved into multiple sources in the MIPS observations and local gas-to-dust ratios were measured to be ∼ 440 and ∼ 250 suggests dust formation and/or significant amounts of ionized gas in these regions. These results support the interpretation that the SMC Tail is a tidal tail recently stripped from the SMC that includes gas, dust, and young stars.

Accepted for publication in ApJ Letters Available from arXiv:0811.2789 and from http://www.stsci.edu/∼kgordon/papers/paper 150.html

The LMC’s Largest Molecular Cloud Complex: Spitzer Analysis of Embedded Star Formation

Remy Indebetouw^{1,2}, Barbara Whitney, Akiko Kawamura, Toshikazu Onishi, Margaret Meixner, Marilyn Meade, Brian Babler, Joe Hora, Karl Gordon, Chad Engelbracht, Miwa Block and Karl Misselt

1University of Virginia, USA 2National Radio Astronomy Observatory, USA

We present a mid-infrared analysis of star-formation activity in the Large Magellanic Cloud’s molecular ridge (south of 30 Doradus). The Magellanic Clouds are a rare laboratory in which extragalactic star-formation diagnostics can be tested at high spatial resolution. The southern part of the molecular ridge is particularly interesting as a potential extreme in the range of molecular cloud conditions, because of the apparent paucity (in optical tracers) of star formation compared to its gas mass. Our Spitzer observations are sensitive to protostars > ∼3 M_{⊙}, and we estimate a total starformation luminosity of 5 × 10^{6} L_{⊙} in 2 × 10^{6} M_{⊙} of molecular material. Detailed modeling of individual infrareddetected star-formation regions yields a total mass of star formation in the region consistent with that predicted by the gas surface density via the Schmidt-Kennicutt relation. The star-formation activity is distributed in rather low-luminosity regions, so the total star-formation rate determined by our infrared analysis is higher than would be predicted simply by the total H i and 24 μm luminosities. Detailed analysis in very nearby galaxies, like the Magellanic Clouds, allows us to test and better understand the scaling relations used in unresolved and distant star-formation regions. Finally, we analyze the star-formation regions in the context of their individual molecular clouds and find that clouds with a higher ratio of CO mass to virial mass are more vigorously forming stars.

Published in AJ, 136, 1442 (2008)

The Optical Gravitational Lensing Experiment. The OGLE-III Catalog of Variable Stars. II. Type II Cepheids and Anomalous Cepheids in the

Large Magellanic Cloud

I. Soszyński^{1}, A. Udalski^{1}, M.K. Szymański^{1}, M. Kubiak^{1}, G. Pietrzyński^{1,2}, Ł. Wyrzykowski^{1,3}, O. Szewczyk^{1,2}, K. Ulaczyk^{1} and R. Poleski^{1}

1Warsaw University Observatory, Al. Ujazdowskie 4, 00-478 Warszawa, Poland 2Universidad de Concepción, Departamento de Física, Casilla 160-C, Concepción, Chile 3Institute of Astronomy, University of Cambridge, Madingley Road, Cambridge CB3 0HA, UK

In the second part of the OGLE-III Catalog of Variable Stars (OIII-CVS) we present 197 type II Cepheids and 82 anomalous Cepheids in the Large Magellanic Cloud (LMC). The sample of type II Cepheids consists of 64 BL Her stars, 96 W Vir stars and 37 RV Tau stars. Anomalous Cepheids are divided into 61 fundamental-mode and 21 first-overtone pulsators. These are the largest samples of such types of variable stars detected anywhere outside the Galaxy. We present the period-luminosity and color–magnitude diagrams of stars in the sample. If the boundary period between BL Her and W Vir stars is adopted at 4 days, both groups differ significantly in (V − I) colors. We identify a group of 16 peculiar W Vir stars with different appearance of the light curves, brighter and bluer than ordinary stars of that type. Four of these peculiar W Vir stars show additional eclipsing modulation superimposed on the pulsation light curves. Four other stars of that type show long-period secondary variations which may be ellipsoidal modulations. It suggests that peculiar W Vir stars may be related to binarity. In total, we identified seven type II Cepheids simultaneously exhibiting eclipsing variations which is a very large fraction compared to classical Cepheids in the LMC. We discuss diagrams showing Fourier parameters of the light curve decomposition against periods. Three sharp features interpretted as an effect of resonances between radial modes are detectable in these diagrams for type II Cepheids.

Accepted for publication in Acta Astronomica Available from arXiv:0811.3636

Properties of WNh stars in the Small Magellanic Cloud: evidence for homogeneous evolution

F. Martins^{1}, D.J. Hillier^{2}, J.C. Bouret^{3}, E. Depagne^{4}, C. Foellmi^{5}, S. Marchenko^{6} and A.F. Moffat^{7}

1GRAAL - CNRS, France 2University of Pittsburgh, USA 3LAM - CNRS, France 4Las Cumbres Observatory, USA 5LAOG - CNRS, France 6University of Western Kentucky, USA 7University of Montréal, Canada

We derive the physical properties of three WNh stars in the SMC to constrain stellar evolution beyond the main sequence at low metallicity and to investigate the metallicity dependence of the clumping properties of massive stars. We compute atmosphere models to derive the stellar and wind properties of the three WNh targets. A FUV/UV/optical/near-infrared analysis gives access to temperatures, luminosities, mass loss rates, terminal velocities and stellar abundances. All stars still have a large hydrogen mass fraction in their atmosphere, and show clear signs of CNO processing in their surface abundances. One of the targets can be accounted for by normal stellar evolution. It is a star with initial mass around 40–50 M_{⊙} in, or close to, the core He burning phase. The other two objects must follow a peculiar evolution, governed by fast rotation. In particular, one object is likely evolving homogeneously due to its position blue-ward of the main sequence and its high H mass fraction. The clumping factor of one star is found to be 0.15 ± 0.05. This is comparable to values found for Galactic Wolf-Rayet stars, indicating that within the uncertainties, the clumping factor does not seem to depend on metallicity.

Accepted for publication in A&A Available from arXiv:0811.3564

Formation of the off-center bar in the LMC from a collision with a dark satellite

Kenji Bekki^{1}

1UNSW, Australia

Recent observations on structural properties of the Large Magellanic Cloud (LMC) based on the Deep Near-Infrared Southern Sky Survey (DENIS) and Two Micron All-Sky Survey (2MASS) have revealed that the LMC has an offcenter bar even in the older stellar populations. Previous dynamical models including tidal interaction between the LMC, the Small Magellanic Cloud (SMC), and the Galaxy, however, did not reproduce so well the older off-center bar. We here show that the off-center bar can be formed if the LMC with an already existing bar can collide with a low-mass Galactic subhalo as massive as 5 × 10^{8} M_{⊙} (corresponding roughly to a few % of the LMC mass). The simulated stellar distribution after the collision appears to show an “off-center bar”, not because the center of the bar significantly deviates from the dynamical center of the LMC, but because the underlying stellar distribution of the disk is significantly asymmetric with respect to the center of the bar. We discuss whether off-center bars observed in Magellanic-type dwarf galaxies can be formed as a result of tidal interaction with low-mass halos with no or little visible matters.

Accepted for publication in MNRAS Letters Available from arXiv:0811.3279

Gemini/GMOS search of massive binaries in the ionizing cluster of 30 Dor

Guillermo Bosch^{1}, Elena Terlevich^{2} and Roberto Terlevich^{2}

1FCAG-IALP, Argentina 2INAOE, México

If binaries are common among massive stars, it will have important consequences for the derivation of fundamental properties like the cluster age, IMF and dynamical mass. Making use of the multiplexing facilities of Gemini Multi Object Spectrograph (GMOS) we were able to investigate the presence of binary stars within the ionising cluster of 30 Doradus. From a seven epochs observing campaign at Gemini South we detect a binary candidate rate of about 50%, which is consistent with an intrinsic 100% binary rate among massive stars. We find that single epoch determinations of the velocity dispersion give values around 30 km s^{−1} . After correcting the global velocity dispersion for the binary orbital motions, the true cluster velocity dispersion is 8.3 km s^{−1}. This value implies a virial mass of about 4.5 × 10^{5} M_{⊙} or 8 percent of the mass calculated using the single epoch value. The binary corrected virial mass estimate is consistent with photometric mass determinations thus suggesting that NGC 2070 is a firm candidate for a future globular cluster.

Accepted for publication in Astronomical Journal Available from arXiv:0811.4748

Conference Papers

Turbulence: A Probe of the Dynamics and Physics of the Magellanic Stream

Itzhak Goldman^{1,2}

1Department of Exact Sciences, Afeka College, Tel Aviv, Israel 2Department of Astronomy and Astrophysics, School of Physics,Tel Aviv University, Tel Aviv, Israel

A recent paper by Stanimirović et al. (2008) presents quit interesting results from H i observations of the Magellanic Stream (MS) tip. The high spatial resolution of the data reveals rich and complex morphological and kinematic structures; notably four coherent H i substreams extending over angular size of ∼ 20^{∘} were found. We use the position–velocity images of Stanimirović et al. (2008) to derive spatial power spectra for the velocity residuals. These, indicate the presence of a large scale turbulence with size comparable to that of the streams themselves. The turbulent velocity on the largest scale is estimated to be ∼ 15 km s^{−1} . Adopting, a distance of 120 kpc, implies a turbulent largest scale of 40 kpc and timescale for decay of ∼ 3 Gyr. For a turbulence with scale that large, the natural energy source is the tidal interaction between the Magellanic Clouds, and between them and the Milky Way galaxy. The shape of derived turbulence spectrum is used here to obtain constraints on the inclination of the streams and on the density of the emitting neutral hydrogen.

Poster contribution, published in IAU Symposium 256, ”The Magellanic System: Stars, Gas, and Galaxies”, Jacco Th. van Loon & Joana M. Oliveira (eds.) Available from arXiv:0809.4523

Properties of X-ray Binaries in the Magellanic Clouds from RXTE and Chandra Observations

R.H.D. Corbet^{1,2}, M.J. Coe^{3}, K.E. McGowan^{3}, M.P.E. Schurch^{3}, L.J. Townsend^{3}, J.L. Galache^{4} and F.E. Marshall^{2}

1University of Maryland, Baltimore County/CRESST, USA 2NASA Goddard Space Flight Center, USA 3School of Physics and Astronomy, University of Southampton, UK 4Harvard-Smithsonian Center for Astrophysics, USA

The X-ray binary population of the SMC is very different from that of the Milky Way consisting, with one exception, entirely of transient pulsating Be/neutron star binaries. We have now been monitoring these SMC X-ray pulsars for over 10 years using the Rossi X-ray Timing Explorer with observations typically every week. The RXTE observations have been complemented with surveys made using the Chandra observatory. The RXTE observations are non-imaging but enable detailed studies of pulsing sources. In contrast, Chandra observations can provide precise source locations and detections of sources at lower flux levels, but do not provide the same timing information or the extended duration light curves that RXTE observations do. We summarize the results of these monitoring programs which provide insights into both the differences between the SMC and the Milky Way, and the details of the accretion processes in X-ray pulsars.

Oral contribution, published in IAU Symposium 256, ”The Magellanic System: Stars, Gas, and Galaxies”, Jacco Th. van Loon & Joana M. Oliveira (eds.), CUP Available from arXiv:0810.4915

Warm gas accretion onto the Galaxy

J. Bland-Hawthorn^{1}

1University of Sydney, Australia

We present evidence that the accretion of warm gas onto the Galaxy today is at least as important as cold gas accretion. For more than a decade, the source of the bright Hα emission (up to 750 mR) along the Magellanic Stream has remained a mystery. We present a hydrodynamical model that explains the known properties of the Hα emission and provides new insights on the lifetime of the Stream clouds. The upstream clouds are gradually disrupted due to their interaction with the hot halo gas. The clouds that follow plough into gas ablated from the upstream clouds, leading to shock ionisation at the leading edges of the downstream clouds. Since the following clouds also experience ablation, and weaker Hα (100–200 mR) is quite extensive, a disruptive cascade must be operating along much of the Stream. In order to light up much of the Stream as observed, it must have a small angle of attack (∼ 20^{∘}) to the halo, and this may already find support in new H i observations. Another prediction is that the Balmer ratio will be substantially enhanced due to the slow shock. We find that the clouds are evolving on timescales of 100–200 Myr, such that the Stream must be replenished by the Magellanic Clouds at a fairly constant rate (> 0.1 M_{⊙} yr^{−1}). The ablated material falls onto the Galaxy as a warm drizzle; diffuse ionized gas at 10^{4} K is an important constituent of galactic accretion. We consider the stability of H i clouds falling towards the Galactic disk and show that most of these must break down into smaller fragments that become partially ionized. The Galactic halo is expected to have huge numbers of smaller neutral and ionized fragments. When the ionized component is accounted for, the rate of gas accretion is ∼ 0.4 M_{⊙} yr^{−1}, roughly twice the rate deduced from H i observations alone.

Oral contribution, published in IAU Symposium 254, ”The Galaxy Disk in Cosmological Context”, Copenhagen (2008) Available from arXiv:0811.2467

Review Papers

The Star Clusters of the Magellanic System

Basílio X. Santiago^{1}

1Instituto de Física, Univ. Federal do Rio Grande do Sul, Porto Alegre, Brazil

More than 50 years have elapsed since the first studies of star clusters in the Magellanic Clouds. The wealth of data accumulated since then has not only revealed a large cluster system, but also a diversified one, filling loci in the age, mass and chemical abundance parameter space which are complementary to Galactic clusters. Catalogs and photometric samples currently available cover most of the cluster mass range. The expectations of relatively long cluster disruption timescales in the Clouds have been confirmed, allowing reliable assessments of the cluster initial mass function and of the cluster formation rate in the Clouds. Due to their proximity to the Galaxy, Magellanic clusters are also well resolved into stars. Analysis of colour–magnitude diagrams (CMDs) of clusters with different ages, masses and metallicities are useful tools to test dynamical effects such as mass loss due to stellar evolution, two-body relaxation, stellar evaporation, cluster interactions and tidal effects. The existence of massive and young Magellanic clusters has provided insight into the physics of cluster formation. The magnitudes and colours of different stellar types are confronted with stellar evolutionary tracks, thus constraining processes such as convective overshooting, stellar mass-loss, rotation and pre main-sequence evolution. Finally, the Magellanic cluster system may contribute with nearby and well studied counterparts of recently proposed types of extragalactic clusters, such as Faint Fuzzies and Diffuse Star Clusters.

Published in IAU Symposium 256, ”The Magellanic System: Stars, Gas, and Galaxies”. Eds. Jacco Th. van Loon & Joana M. Oliveira, CUP Available from arXiv:0810.1514

Abundances of massive stars: some recent developments

T. Morel^{1}

1Institut d’Astrophysique et de Géophysique, Université de Liège, Belgium

Thanks to their usefulness in various fields of astrophysics (e.g., mixing processes in stars, chemical evolution of galaxies), the last few years have witnessed a large increase in the amount of abundance data for early-type stars. Two intriguing results emerging since the last reviews on this topic will be discussed: (a) nearby OB stars exhibit metal abundances generally lower than the solar/meteoritic estimates; (b) evolutionary models of single objects including rotation are largely unsuccessful in explaining the CNO properties of stars in the Galaxy and in the Magellanic clouds.

Published in ”Evolution and Pulsation of Massive Stars on the Main Sequence and Close to it” (Liège, July 2008), Communications in Asteroseismology (invited review) Available from arXiv:0811.4114 and from http://www.astro.ulg.ac.be/∼morel/preprints.html

Job Advert

Côte d’Azur Observatory, Nice, France Post-doctoral position in automatic stellar classification

The Observatory of the Côte d’Azur (Nice, France) calls for applications to a postdoctoral position to work on the automatic classification of the stellar spectra collected by ESO instruments.

The successful applicant will work in close collaboration with P. de Laverny, A. Recio-Blanco, V. Hill and A. Bijaoui, members of the Gaia group of the Côte d’Azur Observatory. Their main scientific interests are stellar evolution and chemical abundances, stellar populations, galactic archeology, automatic classification and the preparation of the ESA/Gaia mission. In this group, a specific algorithm is developed (MATISSE, Recio-Blanco et al., 2006, MNRAS, 370, 461) in order to automatically derive the stellar atmospheric parameters and chemical abundances from spectra collected with current instruments or future ones, such as the Gaia/RVS. In the framework of a collaborative work with ESO, this algorithm will be applied to the ESO stellar spectra in order to provide to the whole astronomical community the stellar parameters together with the reduced spectra directly from the ESO archives.

We are particularly interested by applicants with previous experiences in the field of observational stellar astrophysics, stellar atmospheres, spectroscopic reduction/analysis and/or chemical analysis, but all applicants with experience in related areas will also be considered. Preference will also be given to candidates with strong independent research programs in these areas, as the applicant is expected to carry on his own research in parallel.

The position is for one year renewable, subject to performance and (very likely) extended funding. The starting date has some flexibility within the first term of 2009. Funds for travel and research will be available, as well as easy access to computing facilities.

Applicants are requested to send a CV, a list of publications, and a brief (3 pages) description of past/future research, accomplishments and relevant technical experiences. This material together with three letters of reference should be sent to Patrick de Laverny. The deadline to apply is 10 December 2008, and further information can be directly requested from:

Patrick de Laverny email : laverny@oca.eu Observatoire de la Côte d’Azur BP4229 F-06304 Nice cedex 4